SMOLIRA GOLF CORP.
2017 and 2018 Balance Sheets
Assets Liabilities and Owners’ Equity
2017 2018 2017 2018
  Current assets   Current liabilities
      Cash $ 24,136 $ 25,000       Accounts payable $ 24,084 $ 28,000
      Accounts receivable 13,348 16,100       Notes payable 19,000 11,700
      Inventory 26,722 28,000       Other 12,471 18,300
        Total $ 64,206 $ 69,100         Total $ 55,555 $ 58,000
  Long-term debt $ 78,000 $ 91,532
  Owners’ equity
      Common stock and paid-in surplus $ 60,000 $ 60,000
      Accumulated retained earnings 204,346 223,468
  Fixed assets
  Net plant and equipment $ 333,695 $ 363,900   Total $ 264,346 $ 283,468
  Total assets $ 397,901 $ 433,000   Total liabilities and owners’ equity $ 397,901 $ 433,000
SMOLIRA GOLF CORP.
2018 Income Statement
  Sales $ 383,818
  Cost of goods sold 257,000
  Depreciation 48,300
  Earnings before interest and taxes $ 78,518
  Interest paid 15,200
  Taxable income $ 63,318
  Taxes (24%) 15,196
  Net income $ 48,122
      Dividends $ 29,000
      Retained earnings 19,122

The company's profit margin is ____ percent. (Use year-end figures rather than average values where appropriate and enter your answer as a percent rounded to 2 decimal places).

The total asset turnover is ____ times. (Round your answer to two decimal places).

The equity multiplier is _____ times. (Round your answer to two decimal places).

Using the DuPont Identity, the company's ROE is _____ percent. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places).
Answer #1

a)Net profit margin = Net income /sales

                 = 48122 /383818

                 = .1253771 or 12.53771% (rounded to 12.54%)

b)Total asset turnover = Sales /Total asset

                     = 383818 / 433000

                    =.88642 (rounded to .89 times)

c)Equity Multiplier = Total asset /Total equity

               = 433000/283468

                = 1.52751 (rounded to 1.53 times)

d)Return on equity = Profit margin * Total asset turnover *Equity multiplier

            = .1253771 * .88642 * 1.52751

            = .1698 or 16.98%